OpenStudy (anonymous):

Find dy/dx by implicit differentiation. 2x^3 + x^2y - xy^3 = 1

OpenStudy (amistre64):

d/dx= Dx; dy/dx = y'; and dx/dx=x'=1 Dx(2x^3) = 6x^2(x') Dx(yx^2) = 2xy(x') + y'x^2 Dx(xy^3)= y^3(x') + x 3y^2 y' Dx(1)= 0(x') 6x^2 +2xy + y'x^2 - y^3 - 3xy^2 y' = 0 6x^2 +2xy -y^3 +y'(x^2-3xy^2) = 0 y'(x^2 -3xy^2)=(-6x^2 -2xy +y^3) y'=(-6x^2 -2xy +y^3)/(x^2 -3xy^2) If im lucky, that works out :)

OpenStudy (anonymous):

thank u!